OpenStudy (anonymous):

sqrt135b^2c^3d timessqrt5b^2d

OpenStudy (timaashorty):

\[\sqrt135b^2C ^{3d}* \sqrt 5b ^{2d}\] like that?

OpenStudy (anonymous):

the 135b^2c^3d is all under one sqrt symbol, and the 5b^2d is under another sqrt symbol

OpenStudy (timaashorty):

\[\sqrt{135b^2c^{3d}}* \sqrt {5b^{2d} }\] like this?

OpenStudy (timaashorty):

and what do you need to find?

OpenStudy (anonymous):

I'm sorry, it says simplify and solve the equation

OpenStudy (timaashorty):

do you have answer choices?

OpenStudy (anonymous):

Unfortunately not

OpenStudy (timaashorty):

I would show how to solve it, but I really got to go. But I submitted it in a calculator which will help: http://www.wolframalpha.com/input/?i=sqrt%28135b%5E2c%5E%283d%29%29*+sqrt+%285b%5E%282d%29%29

OpenStudy (timaashorty):

Btw, welcome to openstudy!(:

OpenStudy (anonymous):

Thank you!! I appreciate it very much!!!

OpenStudy (timaashorty):

You're very welcome